# Ion Coulomb crystals: an exotic form of condensed matter

### v1

@@ II.2.3 Normal modes
In a frame that co-rotates with the ion crystal, individual ions in the cyclotron modes undergo an approximate clockwise circular motion about their equilibrium positions. In contrast, the ${\mathbf{E}}\times{\mathbf{B}}$ modes on average exhibit counterclockwise motion [291]. Movies that show the eigenvectors of the planar normal modes of a single plane crystal in a Penning trap can be found in the supplementary material of Ref. [329]. In the rotating frame, both the cyclotron modes and ${\mathbf{E}}\times{\mathbf{B}}$ modes are positive energy modes that have the conventional property that an increase in mode amplitude corresponds to an increase in mode energy. However, the ${\mathbf{E}}\times{\mathbf{B}}$ and cyclotron modes are more complicated than the drumhead modes, which, in the linear approximation, behave like $N$ independent simple harmonic oscillators. In contrast to a simple harmonic oscillator, the time-averaged kinetic and potential energy are not equal for the cyclotron or the ${\mathbf{E}}\times{\mathbf{B}}$ modes. The cyclotron modes are dominated by kinetic energy and the ${\mathbf{E}}\times{\mathbf{B}}$ modes by potential energy. For the normal mode analysis of Fig. 25, which employs conditions relevant for recent experimental work [291] with $^{9}\textrm{Be}^{+}$ and parameters $(\Omega_{c},\omega_{z},\omega_{r})/(2\pi)=(7.60,1.59,0.180)\;\textrm{MHz}$, the ratio of potential to kinetic energy for a typical ${\mathbf{E}}\times{\mathbf{B}}$ mode is of order 200 and while the ratio of potential to kinetic energy for a typical cyclotron mode is 0.005.
@@
The large magnetic field of a Penning trap therefore produces a separation of thermal fluctuations into potential energy fluctuations that are mainly captured by the ${\mathbf{E}}\times{\mathbf{B}}$ branch and kinetic energy fluctuations that are predominantly captured by the cyclotron branch. This separation has consequences for Doppler laser cooling. Doppler laser cooling couples to velocities and kinetic energy, and therefore efficiently cools cyclotron motion. On the other hand, the ${\mathbf{E}}\times{\mathbf{B}}$ modes are dominated by the potential energy associated with positional fluctuations of the ions from their equilibrium positions and are not efficiently cooled by Doppler laser cooling. In addition, recent simulations show that the equilibration rate between the cyclotron branch and the ${\mathbf{E}}\times{\mathbf{B}}$ branch is exponentially weak in the ratio of the frequency of the cyclotron branch to the ${\mathbf{E}}\times{\mathbf{B}}$ branch ($\sim\Omega_{c}/\omega_{m}$[309].
@@
Evidence for elevated ${\mathbf{E}}\times{\mathbf{B}}$ mode temperatures has been obtained from simulations [161] and from the broadening of the spectrum of the drumhead modes, which is sensitive to fluctuations in the ion positions [277, 276, 291]. This motivates developing techniques for improved measurements of the temperature of the in-plane modes, in particular, for the ${\mathbf{E}}\times{\mathbf{B}}$ modes, and generalizing techniques like “axialization”[247], which use time-dependent electric field gradients to couple and equilibrate the cyclotron and magnetron motions of single or a few ions stored in Penning traps. Doppler cooling simulations of single-plane ion crystals in a Penning trap have indicated that efficient cooling of the ${\mathbf{E}}\times{\mathbf{B}}$ modes is possible by overlapping the bandwidths of the ${\mathbf{E}}\times{\mathbf{B}}$ and drumhead modes through appropriate tuning of the ion crystal parameters [161].
@@ II.2.4 Structural instability of the planar geometry
Single-plane crystals are strictly two-dimensional geometries that are stable provided that the transverse confinement is sufficiently strong. By decreasing the transverse trap strength below a critical value, the plane becomes unstable and the crystal undergoes a transition to a three-dimensional structure [94]. This instability, driven by the drumhead modes at the lowest frequency, manifests itself in the formation of a three-plane structure, illustrated in Fig. 26. We call this structure the buckled phase, and the transition into this phase the buckling transition.
The theoretical prediction of the buckling transition of Ref. [94] is based on minimization of the potential energy. The condition for stability of the single-plane structure can be written as $\sigma a_{0}^{2}\lesssim 1.1$, where $a_{0}$ is the characteristic length emerging from the competition of Coulomb repulsion and the transverse harmonic confinement. Equivalently, stability requires the trap frequency $\omega_{z}$ to exceed the value $\omega_{z}^{\rm MF}$, with
@@
The role of ion kinetic energy in the buckling transition has been examined in a field theory analysis presented in Ref. [250]. Due to their large mass, ions typically have negligible kinetic energy compared to their potential energy. However, at the buckling transition, where the Coulomb and confining potential energies balance each other, kinetic energy becomes disproportionately significant. This leads to thermal and quantum fluctuations that alter the nature of the buckling instability. Notably, the transition from one to three planes occurs via a two-stage process, where the planar (disordered) and buckled (ordered) phases are separated by an intermediate, partially ordered phase. As discussed below, this behavior is captured by mapping the field theory onto a six-state clock model, a well-studied framework in statistical physics [163, 240, 198].
Figure 26: A single-plane, triangular crystal undergoes a so-called buckling transition to three planes, each with triangular structure. Left panel: Height pattern in the ordered phase. The triangular lattice is split into 3 sublattices, one in the $z=0$ plane (gray circles), one above it (red), and one below it (blue). There are $3!=6$ inequivalent configurations, corresponding to the possible height assignments to the sublattices. Right panel: The first Brillouin zone of the single-plane crystal is shown, together with the wave vectors $\bf K$} and ${\bf K}^{*}=-{\bf K}$ of the height pattern. From 250
The first step in the field theory analysis is to identify the order parameter of the buckled phase. The pattern shown in Fig. 26, for example, is encoded by the expression
@@
This naturally suggests a mapping to the six-state clock model [163, 240, 198]. The mapping is performed for transverse frequencies $\omega_{z}$ near the mean-field critical value $\omega_{z}^{\rm MF}$ (Eq. (58)). In this regime, the displacements $z_{i}$, and consequently $|\psi|$, remain much smaller than the in-plane interparticle distance, allowing for corrections to the planar Coulomb interactions to be incorporated via a multipole expansion. By assuming that the complex-valued order parameter $\psi(\bm{\rho},t)$ varies slowly in space and time, a gradient expansion can be employed by treating the discrete planar distribution as a continuum. This leads to a Ginzburg-Landau (GL) theory expressed in terms of $\psi(\bm{\rho},t)$, whose form is constrained by the symmetries of the triangular lattice. The explicit formulation of the GL model can be found in Ref. [250]. Here, we present its form when the transverse trap frequency $\omega_{z}$ falls below the mean-field threshold $\omega_{z}^{\rm MF}$. In this regime, the order parameter $\psi$ acquires a finite amplitude, $|\psi|=|\psi_{0}|$, which is effectively described by the mean-field model. Meanwhile, the phase dynamics of $\theta(\bm{\rho},t)$ are governed by the Lagrangian density:
${\mathcal{L}}_{\rm GL}=\frac{\rho_{s}^{0}}{2}\left(\frac{1}{c^{2}}|\partial_{t}\theta|^{2}-|\nabla\theta|^{2}\right)-h_{6}\cos(6\theta)$ (60)
@@
Figure 27: Phase diagram of the structural instability of a planar ion crystal in the parameter space determined by the temperature $T$ and $r=\omega_{z}-\omega_{z}^{\rm MF}$. For $r>0$ the ions form a stable, single-plane crystal. At $r<0$, in the white region, the transverse structure is disordered because of either thermal (for $T>0$) or quantum ($T=0$) fluctuations. The external line at $T_{\mathrm{KT}}$ separates the disordered phase from a critical phase, where the ions are distributed along three layers and exhibit quasi long-range order in the transverse direction. The second line at $T_{6}$ indicates the onset of the long-range ordered buckled phase, where the ions arrange in three static planes as illustrated in Fig. 26. The value $r_{c}<0$ indicates the critical point at $T=0$. The dashed line at $T_{6}^{*}$ is a crossover separating quantum and classical behaviors within the ordered phase. From 250.
Figure 27 presents the phase diagram as a function of two control parameters: temperature $T$ and the transverse trap frequency, here expressed as the difference $r=\omega_{z}-\omega_{z}^{\rm MF}$ from the mean-field value $\omega_{z}^{\rm MF}$. At finite temperatures, the model predicts a two-stage thermal transition. At high temperatures, for $T>T_{\mathrm{KT}}$, the system remains in a fully disordered phase, where ions undergo thermal fluctuations around the single-plane structure. In this phase, correlation functions of the atoms’ heights decay exponentially with distance and time. At low temperatures, below a critical temperature denoted by $T_{6}$, the system enters a fully ordered phase, where the ions’ equilibrium positions form a static, three-plane buckled structure. Both transitions are of the Kosterlitz-Thouless (KT) type, with critical temperatures determined by the conditions
@@
As the temperature $T$ decreases, quantum fluctuations become increasingly dominant, ultimately driving a phase transition at a single quantum critical point $(T=0,\,r=r_{c})$ [240, 198]. For temperatures above $T_{6}(r)$, the clock term $\cos(6\theta)$ in Eq. (60) is irrelevant in the renormalization group sense. Consequently, the model falls within the universality class of the $XY$ model in $D=2+1$ dimensions, implying that the stiffness $\rho_{s}$ scales as $r$ approaches $r_{c}$ according to
@@
II.3 Three dimensional crystals
Three-dimensional (3D) trapped ion crystals have been prepared in Penning and rf ion traps since the early days of laser cooling [78, 332, 123]. While 1D and 2D ion crystals have been limited to $\sim 500$ ions, 3D crystals can reach sizes greater than $10^{5}$ ions. Large, 3D crystals provide intriguing opportunities for developing new control and characterization techniques as well as challenges. In fact, the control and characterization techniques employed with 3D crystals is not as advanced as that developed for 1D and 2D crystals. To date, only Doppler cooling has been applied with 3D crystals. We review here the types of 3D crystal configurations that have been prepared in the lab, discuss the importance of finite size effects, and summarize the long wavelength normal modes of 3D crystals.
@@ II.3.1 Ground state configuration
Before discussing the detailed structure of three-dimensional ion Coulomb crystals, it is useful to consider the cold fluid plasma limit of a single species ion crystal, in which the ions are approximated by a continuous charge distribution in space. This is also called the zero-temperature, mean-field plasma [98]. For simplicity, consider an effective harmonic confinement potential in all three dimensions, which is routinely achieved in Penning and quadrupole rf traps. At low temperatures this gives rise to constant density ion crystals with sharp boundaries. In the cold fluid limit, the continuous charge distribution is uniform in space with approximately the same boundary as the ion crystal. The energy required to assemble the continuous charge distribution does not scale extensively with (that is, it is not proportional to) the number of trapped charges.
The cold fluid limit is useful for determining the overall crystal shape but the specific lattice structure is dictated by the discreteness of the charges and, more precisely, by the short distance contributions to the energy. The latter is the so-called correlation energy and scales extensively with the lattice size. Three-dimensional strongly coupled OCPs are peculiar systems in that three different lattice configurations have nearly the same correlation energy per ion in the infinite (bulk) lattice and at zero temperatures. These configurations are body centered cubic (bcc), face centered cubic (fcc) and hexagonal close packed (hcp). The bcc lattice is the zero temperature ground state [41, 153], but the bulk energy per ion in the fcc and hcp are less than $1\times 10^{-4}$ higher than the one in the bcc configuration [9, 91].
@@
where $\bm{r}$ denotes the spherical radius vector. Due to the spherical symmetry of the confining potential, the global structure of finite Coulomb crystals is expected to be spherical as well. Crystals containing up to a few thousand ions minimize their potential energy as the temperature goes to zero by organizing themselves in concentric spherical shells with a specific maximum “magic” number of ions in each shell [133, 90]; see Fig. 28. Within each shell the ions furthermore form a two-dimensional triangular lattice (like that for single-plane crystals; see Sec. II.2.2), but with a number of defects to account for the curvature of the surface (see Fig. 28). The ground state (zero temperature) configurations are traditionally found by Monte Carlo methods or Molecular Dynamics (MD) simulations combined with careful annealing procedures. It is challenging, however, to determine the absolute ground state configurations already for a modest number of ions ($\sim$100), since, as is well-known from cluster physics, the number of slightly excited metastable configurations increases exponentially with the particle number. Combined with the small deviations from a perfect harmonic confining potential in real experiments, predictions of the exact structures obtainable experimentally become extremely challenging.
@@
The uncertainty in the predicted magic numbers of ions within each shell is, though, so small that when comparing a given observed spherical shell structure for crystals consisting of up to a $\sim$ thousand ions in the laboratory, one typically obtains an estimate of the number of ions at an uncertainty level much better than the square-root of the number of ions. Hence, structural information represents a reliable way to obtain information on the number of ions in a real crystal.
@@
Figure 28: Molecular dynamics simulations (MD) of spherical symmetric confined ion Coulomb crystals. a) Observed shell structure as a function of the number of ions $N$ in the simulations. When a sign ++ appears in the column StructureStructure, it means a new shell starts forming, and the number NN denotes the corresponding ”magic” number. In all simulations the final Coulomb coupling parameter $\Gamma$ is larger than 105. b) Illustration of the position of the ions in the outer shell of a 5000 ion system. Note the general hexagonal structure with a few point defects, such as the one below the center. The length scale is the Wigner-Seitz radius $a_{AW}$ From 133.
@@
Ellipsoidal and spheroidal geometries— A similar shell structure is obtained for the more general ellipsoidal ion crystal shapes (or spheroidal if the confinement in two orthogonal directions is equal). The Coulomb potential energy of the crystal is minimized by forming a shell on the boundary of the crystal that conforms with the overall crystal shape determined by the confining potential. This shell structure propagates into the interior of the ion crystal through the formation of commensurately shaped shells of smaller size. Experimentally, shell structure is routinely observed in trapped 3D ion crystals [123, 84, 146, 281]. Figure 29 shows some experimental observations of ion crystal shell structure.
[…]
@@
An interesting question is at what size will it become energetically favorable for an ion crystal to have a bcc lattice as its ground state configuration. Two different theoretical studies based on molecular dynamics simulations with spherical OCPs and ion numbers ranging from a few thousand to greater than $10^{5}$ ions indicate that when $N>10^{4}$ the configuration consisting of a bcc lattice with relaxed surfaces has lower energy than shell structure [319, 135]. Interestingly, if the molecular dynamics simulation started with randomly distributed ions and was slowly annealed, the formation of spherical shells that advanced from the boundary of the crystal to the center was always observed. On the other hand, starting with a spherical cutout of a bcc lattice and annealing resulted in a bcc crystal in the interior surrounded by a few shells at the boundary. For $N>10^{4}$ the spherical bcc crystal had the lowest energy.
@@
Experimentally, stable bcc crystals have been observed in the interior of large ion crystals formed in Penning traps through Bragg scattering of the laser-induced resonance fluorescence as well as through direct imaging of the ion resonance fluorescence [29]. Long range 3D periodic order was observed in approximately spherical crystals with as few as $N\approx 5\times 10^{4}$ [308]. The crystal diameter was $\sim 70a_{\mathrm{WS}}$ where $a_{\mathrm{WS}}$ is the Wigner-Seitz radius defined in Eq. (2). With $N>2\times 10^{5}$ ions (crystal diameter $130a_{\mathrm{WS}}$), bcc crystals were the only type of 3D periodic crystal that was reported [155]. This work could only place a crude lower limit on the observed bcc crystal size of $20a_{\mathrm{WS}}$, which is a small fraction of the entire ion crystal size.
@@
Cylindrical geometry— A different well-characterized trapping situation for three-dimensional crystals is one where ions are confined by an isotropic harmonic potential in two dimensions while being free to move along the third dimension. Here, the parameter governing the crystal structure is the linear density of charged particles along the unconfined axis. For low linear densities concentric cylindrical shell structures with varying surface structures within the shells represent the minimum energy states, with eventually a core which will contain translational lattice structures as the linear density increases (See Fig. 30).
@@
Figure 30: Molecular dynamics simulation of cross sections of cylindrically confined Coulomb crystals for various normalized linear charge densities $\lambda$. For the lowest charge density, $\lambda$=0.266, the charged particles form a one-dimensional string (a single dot in the cross section view), whereas, as the charge density increases, an increasing number of concentric shells, sometimes with a central string, form. The figure is adapted from 136, where more details on the simulations can be found.
@@
In the laboratory this idealized situation cannot be fully implemented. It can effectively be realized by choosing a ring shaped two-dimensional rf quadrupole trap, as has been pioneered by Prof. H. Walther at MPQ [328, 22] (see Fig. 31). As long as the ring’s circumference is much larger than the typical ion-ion distance and the radial extent of the crystal, only very small deviations from the ideal infinite situation are expected.
@@
Figure 31: Sketch of an rf ring trap experiment used to study ion Coulomb crystals under effective cylindrical confinement (a), together with a few examples of fluorescence images of ion crystalline structures obtained with different linear densities (b). The observed helical structures are are sketched below the images. From 22.

### v2

@@ II.2.3 Normal modes
In a frame that co-rotates with the ion crystal, individual ions in the cyclotron modes undergo an approximate clockwise circular motion about their equilibrium positions. In contrast, the ${\mathbf{E}}\times{\mathbf{B}}$ modes on average exhibit counterclockwise motion [326]. Movies that show the eigenvectors of the planar normal modes of a single-plane crystal in a Penning trap can be found in the supplementary material of Ref. [365]. In the rotating frame, both the cyclotron modes and ${\mathbf{E}}\times{\mathbf{B}}$ modes are positive-energy modes. They exhibit the conventional property that an increase in mode amplitude corresponds to an increase in mode energy. However, the ${\mathbf{E}}\times{\mathbf{B}}$ and cyclotron modes are more complicated than the drumhead modes, which, in the linear approximation, behave like $N$ independent simple harmonic oscillators. In contrast to a simple harmonic oscillator, the time-averaged kinetic and potential energy are not equal for the cyclotron or the ${\mathbf{E}}\times{\mathbf{B}}$ modes. The cyclotron modes are dominated by kinetic energy and the ${\mathbf{E}}\times{\mathbf{B}}$ modes by potential energy. For the normal mode analysis of Fig. 25, which employs conditions relevant for recent experimental work [326] with $^{9}\textrm{Be}^{+}$ and parameters $(\Omega_{c},\omega_{z},\omega_{r})/(2\pi)=(7.60,1.59,0.180)\;\textrm{MHz}$, the ratio of potential to kinetic energy for a typical ${\mathbf{E}}\times{\mathbf{B}}$ mode is of order 200 and while the ratio of potential to kinetic energy for a typical cyclotron mode is 0.005.
@@
The large magnetic field of a Penning trap therefore produces a separation of thermal fluctuations into potential energy fluctuations that are mainly captured by the ${\mathbf{E}}\times{\mathbf{B}}$ branch and kinetic energy fluctuations that are predominantly captured by the cyclotron branch. This separation has consequences for Doppler laser cooling. In fact, Doppler laser cooling exploits the dependence of the scattering rate on the velocity via the Doppler shift, and therefore efficiently cools cyclotron motion. On the other hand, the ${\mathbf{E}}\times{\mathbf{B}}$ modes are dominated by the potential energy associated with positional fluctuations of the ions from their equilibrium positions and are therefore inefficiently laser-cooled. In addition, recent simulations show that the equilibration rate between the cyclotron branch and the ${\mathbf{E}}\times{\mathbf{B}}$ branch is exponentially weak in the ratio of the frequency of the cyclotron branch to the ${\mathbf{E}}\times{\mathbf{B}}$ branch ($\sim\Omega_{c}/\omega_{m}$[344].
@@
Evidence for elevated temperatures of the ${\mathbf{E}}\times{\mathbf{B}}$ mode has been obtained from simulations [174] and from the measurement of the broadening of the spectrum of the drumhead modes, which is sensitive to fluctuations in the ion positions [310, 309, 326]. This motivates developing techniques for improved measurements of the temperature of the in-plane modes, in particular, for the ${\mathbf{E}}\times{\mathbf{B}}$ modes, and generalizing techniques like “axialization”[274], which use time-dependent electric field gradients to couple and equilibrate the cyclotron and magnetron motions of single or a few ions stored in Penning traps. Doppler cooling simulations of single-plane ion crystals in a Penning trap have indicated that efficient cooling of the ${\mathbf{E}}\times{\mathbf{B}}$ modes is possible by overlapping the bandwidths of the ${\mathbf{E}}\times{\mathbf{B}}$ and drumhead modes through appropriate tuning of the ion crystal parameters [174].
@@ II.2.4 Structural instability of the planar geometry
Single-plane crystals are strictly two-dimensional geometries that are stable provided that the transverse confinement is sufficiently strong. By decreasing the transverse trap strength below a critical value, the plane becomes unstable and the crystal undergoes a transition to a three-dimensional structure [100]. This instability is driven by the drumhead modes at the lowest frequency and manifests itself in the formation of a three-plane structure, illustrated in Fig. 26. We call this structure the buckled phase, and the transition into this phase the buckling transition.
The theoretical prediction of the buckling transition of Ref. [100] is based on minimization of the potential energy. The condition for stability of the single-plane structure can be written as $\sigma a_{0}^{2}\lesssim 1.1$, where $a_{0}$ is the characteristic length emerging from the competition of the Coulomb repulsion and the transverse harmonic confinement. Equivalently, stability requires the trap frequency $\omega_{z}$ to exceed the value $\omega_{z}^{\rm MF}$, with
@@
Due to their large mass, ions typically have negligible kinetic energy compared to their potential energy. However, at the buckling transition, where the Coulomb and confining potential energies balance each other, kinetic energy becomes disproportionately significant. This leads to thermal and quantum fluctuations that alter the nature of the buckling instability. The role of ion kinetic energy in the buckling transition has been examined in a field theory analysis presented in Ref. [278]. This theory predicts that the transition from one to three planes occurs via a two-stage process, where the planar (disordered) and buckled (ordered) phases are separated by an intermediate, partially ordered phase. This behavior is captured by mapping the field theory onto a six-state clock model, a well-studied framework in statistical physics [176, 267, 218]. Below we review the basic steps leading to the mapping and its predictions.
Figure 26: (Color online) A single-plane, triangular crystal undergoes a so-called buckling transition to three planes, each with triangular structure. Left panel: Height pattern in the ordered phase. The triangular lattice is split into 3 sublattices, one in the $z=0$ plane (gray medium circles), one above it (large red circles), and one below it (small blue circles). There are $3!=6$ inequivalent configurations, corresponding to the possible height assignments to the sublattices. Right panel: The first Brillouin zone of the single-plane crystal is shown, together with the wave vectors $\bf K$} and ${\bf K}^{*}=-{\bf K}$ of the height pattern. From 278
The first step of the field theory analysis is to identify the order parameter of the buckled phase. The pattern shown in Fig. 26, for example, is encoded by the expression
@@
This naturally suggests a mapping to the six-state clock model [176, 267, 218]. The mapping is performed for transverse frequencies $\omega_{z}$ near the mean-field critical value $\omega_{z}^{\rm MF}$ (Eq. (58)). In this regime, the displacements $z_{i}$, and consequently $|\psi|$, remain much smaller than the in-plane interparticle distance, which allows corrections to the planar Coulomb interaction to be incorporated via a multipole expansion. By assuming that the complex-valued order parameter $\psi(\bm{\rho},t)$ varies slowly in space and time, a gradient expansion can be employed by treating the discrete planar distribution as a continuum. This leads to a Ginzburg-Landau (GL) theory expressed in terms of $\psi(\bm{\rho},t)$, whose form is constrained by the symmetries of the triangular lattice. The explicit formulation of the GL model can be found in Ref. [278]. Here, we present its form when the transverse trap frequency $\omega_{z}$ falls below the mean-field threshold $\omega_{z}^{\rm MF}$. In this regime, the order parameter $\psi$ acquires a finite amplitude, $|\psi|=|\psi_{0}|$, which is effectively described by the mean-field model. Meanwhile, the phase dynamics of $\theta(\bm{\rho},t)$ are governed by the Lagrangian density:
${\mathcal{L}}_{\rm GL}=\frac{\rho_{s}^{0}}{2}\left(\frac{1}{c^{2}}|\partial_{t}\theta|^{2}-|\nabla\theta|^{2}\right)-h_{6}\cos(6\theta)\,$, (60)
@@
Figure 27: Phase diagram of the structural instability of a planar ion crystal in the parameter space determined by the temperature $T$ and $r=\omega_{z}-\omega_{z}^{\rm MF}$. For $r>0$ the ions form a stable, single-plane crystal. At $r<0$, in the white region, the transverse structure is disordered because of either thermal (for $T>0$) or quantum ($T=0$) fluctuations. The external line at $T_{\mathrm{KT}}$ separates the disordered phase from a critical phase, where the ions are distributed along three layers and exhibit quasi long-range order in the transverse direction. The second line at $T_{6}$ indicates the onset of the long-range ordered buckled phase, where the ions arrange in three static planes as illustrated in Fig. 26. The value $r_{c}<0$ indicates the critical point at $T=0$. The dashed line at $T_{6}^{*}$ is a crossover separating quantum and classical behaviors within the ordered phase. From 278.
Figure 27 illustrates the phase diagram as a function of two control parameters: temperature $T$ and the transverse trap frequency, here expressed as the difference $r=\omega_{z}-\omega_{z}^{\rm MF}$ from the mean-field value $\omega_{z}^{\rm MF}$. At finite temperatures, the model predicts a two-stage thermal transition. At high temperatures, for $T>T_{\mathrm{KT}}$, the system remains in a fully disordered phase, where ions undergo thermal fluctuations around the single-plane structure. In this phase, correlation functions of the atoms’ heights decay exponentially with distance and time. At low temperatures, below a critical temperature denoted by $T_{6}$, the system enters a fully ordered phase, where the ions’ equilibrium positions form a static, three-plane buckled structure. Both transitions are of the Kosterlitz-Thouless (KT) type, with critical temperatures determined by the conditions
@@
As the temperature $T$ decreases, quantum fluctuations become increasingly dominant, ultimately driving a phase transition at a single quantum critical point $(T=0,\,r=r_{c})$ [267, 218]. For temperatures above $T_{6}(r)$, the clock term $\cos(6\theta)$ in Eq. (60) is effectively suppressed by fluctuations. Consequently, the model describing the transition corresponds to the quantum XY model in two space dimensions, implying that the stiffness $\rho_{s}$ scales as $r$ approaches $r_{c}$ according to
@@
II.3 Three-dimensional crystals
Three-dimensional (3D) trapped ion crystals have been prepared in Penning and rf ion traps since the early days of laser cooling [84, 368, 133]. While 1D and 2D ion crystals have been limited to $\sim 500$ ions, 3D crystals can reach sizes greater than $10^{5}$ ions [343]. Large, 3D crystals provide intriguing opportunities for developing new control and characterization techniques because the current techniques employed with 3D crystals is not as advanced as that developed for 1D and 2D crystals. To date, only Doppler cooling has been applied with 3D crystals. We review here the types of 3D crystal configurations that have been prepared in the lab, discuss the importance of finite size effects, and summarize characteristic properties of the vibrational motion.
@@ II.3.1 Ground state configuration
Before discussing the detailed structure of three-dimensional ion Coulomb crystals, it is useful to consider the cold fluid plasma model of a single species ion crystal, in which the ions are approximated by a continuous charge distribution in space. This is also called the zero-temperature, mean-field plasma [105]. For simplicity, consider an effective harmonic confinement potential in all three dimensions, which is routinely achieved in Penning and quadrupole rf traps. At low temperatures this gives rise to ion crystals with constant density and sharp boundaries. In the cold fluid model, the continuous charge distribution is uniform in space with approximately the same boundary as the ion crystal. The energy required to assemble the continuous charge distribution does not scale extensively with (that is, it is not proportional to) the number of trapped charges.
The cold fluid model is useful for determining the overall crystal shape and the number of trapped ions through the volume enclosed by the crystal shape, but the specific lattice structure is dictated by the discreteness of the charges and, more precisely, by the short distance contributions to the energy. The latter is the so-called correlation energy and scales extensively with the lattice size. Three-dimensional strongly coupled OCPs are peculiar systems in that three different lattice configurations have nearly the same correlation energy per ion in the infinite (bulk) lattice and at zero temperatures. These configurations are body centered cubic (bcc), face centered cubic (fcc) and hexagonal close packed (hcp). The bcc lattice is the zero temperature ground state [43, 165], but the bulk energy per ion in the fcc and hcp are less than $1\times 10^{-4}$ higher than the one in the bcc configuration [97, 10].
@@
with $\bm{r}$ the spherical radius vector. Due to the spherical symmetry of the confining potential, the global structure of finite Coulomb crystals is expected to be spherical as well. Crystals containing up to a few thousand ions minimize their potential energy as the temperature goes to zero by organizing themselves in concentric spherical shells with a specific maximum “magic” number of ions in each shell [104, 148, 312, 48]; see Fig. 28. Within each shell the ions form a two-dimensional triangular lattice (like single-plane crystals; see Sec. II.2.2), but with a number of defects to account for the curvature of the surface (see Fig. 28). The ground state (zero temperature) configurations are traditionally found by Monte Carlo methods or molecular dynamics simulations combined with careful annealing procedures. It is challenging, however, to determine the absolute ground state configurations already for a modest number of ions ($\sim$100), since, as is well-known from cluster physics, the number of slightly excited metastable configurations increases exponentially with the particle number. Combined with the small deviations from a perfect harmonic confining potential in real experiments, predictions of the exact structures obtainable experimentally become extremely challenging.
@@
The uncertainty in the predicted magic numbers of ions within each shell is, though, so small that when observing a given spherical shell structure for crystals consisting of up to thousand ions in the laboratory, one can typically obtain an estimate of the number of ions at an uncertainty level much better than the square-root of the number of ions. Hence, measurements of the crystal structure provide a reliable way to obtain information on the number of ions in a real crystal.
@@
Figure 28: Molecular dynamics simulations of spherical symmetric confined ion Coulomb crystals. a) Shell structure as a function of the number of ions $N$. When a sign ++ appears in the column StructureStructure, it means a new shell starts forming. In this case, NN is the corresponding ”magic” number.” In all simulations the final Coulomb coupling parameter $\Gamma$ is larger than 105. b) Illustration of the position of the ions in the outer shell of a 5000 ion system. Note the general hexagonal structure with a few point defects, such as the one below the center. The length scale is the Wigner-Seitz radius $a_{\rm WS}$. From 148.
@@
Ellipsoidal and spheroidal geometries— A similar shell structure is obtained for the more general ellipsoidal ion crystal shapes (or spheroidal if the confinement in two orthogonal directions is equal  [^3] A spheroid is the shape obtained by rotating an ellipse about one of its principal axes). The Coulomb potential energy of the crystal is minimized by forming a shell on the boundary of the crystal that conforms with the overall crystal shape determined by the confining potential. This shell structure propagates into the interior of the ion crystal through the formation of commensurately shaped shells of smaller size. Experimentally, shell structures are routinely observed in trapped 3D ion crystals [90, 133, 158, 315]. See Fig. 29 for experimental images.
Figure 29: Images of the ion fluorescence from trapped-ion crystals exhibiting shell structure. (a) Horizontal and (b) vertical view of an ion crystal consisting of about 2000 Be+ ions stored in a linear rf trap. From 315. (c) Image of the ion fluorescence from three laser beams directed through the center of a $N\approx 15000$ Be+ ion crystal that exhibits shell structure. The ions are stored in a Penning trap and the image, averaged over many rotations of the ion crystal, is obtained along the magnetic field or rotation axis of the ion crystal. The crystal diameter is 0.6 mm. From 133.
@@
An interesting question is at what size will it become energetically favorable for an ion crystal to have a bcc lattice as its ground state configuration. Molecular dynamics simulations with spherical OCPs indicate that the spherical bcc crystal has lower energy than shell structure for $N>10^{4}$. The simulations were performed independently in [354, 147] with ion numbers ranging from a few thousand to greater than $10^{5}$ ions. Interestingly, if the molecular dynamics simulation started with randomly distributed ions and was slowly annealed, the formation of spherical shells was always observed that advanced from the boundary of the crystal to the center. On the other hand, starting with a spherical cutout of a bcc lattice and then annealing resulted in a bcc crystal in the interior surrounded by a few shells at the boundary.
@@
Experimentally, stable bcc crystals have been observed in the interior of large ion crystals formed in Penning traps through Bragg scattering of the laser-induced resonance fluorescence as well as through direct imaging of the ion resonance fluorescence [32]. Long range 3D periodic order was observed in approximately spherical crystals with as few as $N\approx 5\times 10^{4}$ [343]. The crystal diameter was $\sim 70a_{\mathrm{WS}}$ where $a_{\mathrm{WS}}$ is the Wigner-Seitz radius defined in Eq. (2). With $N>2\times 10^{5}$ ions (crystal diameter $130a_{\mathrm{WS}}$), bcc crystals were the only type of 3D periodic crystal that was reported [168]. This work could only place a crude lower limit on the observed bcc crystal size of $20a_{\mathrm{WS}}$, which is a small fraction of the entire ion crystal size. In linear rf traps, stable fcc and bcc crystalline structures have been also observed in Coulomb crystals with more than $10^{4}$ ions through fluorescence imaging [253].
@@
Cylindrical geometry— A different trapping situation for three-dimensional crystals is realized when the ions are confined by an isotropic harmonic potential in two dimensions while being free to move along the third dimension. Here, the parameter governing the crystal structure is the linear density of charged particles along the unconfined axis. For low linear densities the minimum energy states are concentric cylindrical shell structures with varying surface structures within the shells. As the linear density increases the core forms lattice structures that are translationally invariant along the axis (See Fig. 30).
@@
Figure 30: Molecular dynamics simulation of cross sections of cylindrically confined Coulomb crystals for various normalized linear charge densities $\lambda$. For the lowest charge density, $\lambda$=0.266, the charged particles form a one-dimensional string (a single dot in the cross section view), whereas, as the charge density increases, an increasing number of concentric shells, sometimes with a central string, form. The figure is adapted from 145, where more details on the simulations can be found.
@@
In the laboratory this idealized situation cannot be fully implemented. It can effectively be realized by choosing a ring shaped two-dimensional rf quadrupole trap, as has been pioneered by Herbert Walther and collaborators  [25, 364] (see Fig. 31). As long as the ring’s circumference is much larger than the typical ion-ion distance and the radial extent of the crystal, only very small deviations from the ideal infinite situation are expected.
@@
Figure 31: Sketch of the rf ring trap experiment used to study ion Coulomb crystals under effective cylindrical confinement (a), together with some examples of fluorescence images of ion crystalline structures (here helical structures) obtained with different linear densities (b). The helical structures are sketched below the images. See also Fig. 13. From 25.